# Echo State Networks for Spatio-Temporal Area-Level Data

### v1

@@ 3.1 Areal Random Representation
In adddition to the reservoir, the input signal is another important component of the ESN. It can include covariates that account for confounding effects not captured by the time series alone. When additional covariate variables are unavailable, lagged values are typically used as input signals, as their trajectories describe the temporal dynamics of the series (McDermott and Wikle,, 2017, 2019; Yoo and Wikle,, 2023). In spatial statistics, spatial input embedding is widely used to capture the spatial structure found in the data. For example, Daw and Wikle, (2023) use random Fourier features to approximate the stationary Gaussian process kernel. However, there is no inherent way of incorporating neighborhood structures into the ESN. The common approach for capturing spatial features in ESNs involves modeling the EOFs coefficients asscociated with a spatio-temporal series (McDermott and Wikle,, 2017, 2019; Yoo and Wikle,, 2023). This approach, however, requires deciding how many EOFs to include and may lead to a potential loss in forecasting performance, as demonstrated in our data analysis.
We aim to capture spatial dynamics at the input stage, which offers two key advantages. First, manipulating the input stage preserves the desirable properties of the ESN, such as the echo state property, which relies on specific characteristics of the reservoir weights. Second, handling spatial dynamics at the input stage introduces minimal computational overhead, as it avoids being recursively computed at every time step. As mentioned in Section 1, we leverage random representations to capture spatial dynamics at the input stage (Bengio et al.,, 2013; Gallant,, 1987; Sutton et al.,, 2014). Building on the approach of Gallant, (1987), we represent spatial features found in areal data through a dense embedding of input signals, using a structure inspired by SGC layers (Wu et al.,, 2019). We refer to this random representation as the areal random representation. Let $\bm{X}\in\mathbb{R}^{n_{s}\times n_{x}}$ denote the input signal, $\bm{A}\in\mathbb{R}^{n_{s}\times n_{s}}$ the proximity matrix, and $\widetilde{\bm{A}}=\bm{I}+\bm{A}$ the proximity matrix with self-loops included. Let $\widetilde{\bm{D}}$ represent the degree matrix of $\widetilde{\bm{A}}$. The areal random representation is defined as follows:
$\bm{S}$ $=$ $\widetilde{\bm{D}}^{-1/2}\widetilde{\bm{A}}\widetilde{\bm{D}}^{-1/2},$
$\bm{z}^{(k)}$ $=$ $\bm{S}(\bm{U}^{(k)}\odot\bm{X}\cdot\bm{1}),$
@@
where $\odot$ is the Hadamard product, $\bm{U}^{(k)}$, for $k=1,\dots,K$, are $K$ randomly sampled weights of dimension $n_{s}\times n_{x}$, and $\bm{S}$ is the symmetric normalized adjacency matrix as used in GCN (Kipf and Welling,, 2016). Similar to GCN, this embedding method applies the graph Fourier transform to the input signals, transforming them into the graph spectral domain. However, unlike traditional GCN, which uses shared and learned weights, our method employs weights that are randomly sampled from a uniform distribution with a scale parameter $a_{u}$. Furthermore, these weights are sampled independently for each entry in $\bm{X}$, rather than being shared across the entire graph. This design choice is inspired by Liu et al., (2020), who show that non-shared weights can outperform shared weights in GCNs. Additionally, because our weights are not trained, sharing them does not help in preventing overfitting. To further expand the capacity of representation, we generate $K$ different copies of locally averaged random signals for each $\bm{U}^{(k)}$. The resulting embedding matrix $\mathbf{Z}$ is the collection of $\{\mathbf{z}^{(1)},\dots,\mathbf{z}^{(K)}\}$, with dimensions $n_{s}\times K$. Thus, we expand the input signals at each location from dimension $n_{x}$ to $K$.
We apply the areal random representation to the input signals $\widetilde{\bm{x}}_{t}$ at each time step, resulting in the embedded signals $\bm{z}_{t}^{(k)}$, as shown in (5). To integrate these embedded signals into the ESN, we feed the vectorized signals $\widetilde{\bm{z}}_{t}$ as the input to the ESN, as in (4). This implies that when a signal at any location is activated in the reservoir stage, all signals in its neighborhood are also activated. So,
@@
$\bm{h}_{t}$ $=(1-\alpha)\bm{h}_{t-1}+\alpha\widetilde{\bm{h}}_{t}$
$\widetilde{\bm{h}}_{t}$ $=g_{h}(\frac{\nu}{|\lambda_{w}|}\bm{W}_{res}\bm{h}_{t-1}+\bm{W}_{in}\widetilde{\bm{z}}_{t})$
@@
$\bm{z}_{t}^{(k)}$ $=\bm{S}(\bm{U}^{(k)}\odot\widetilde{\bm{x}}_{t}\cdot\bm{1})$ (5)
$\bm{U}^{(k)}$ $\sim\text{Unif}(-a_{u},a_{u}),$
where $\bm{Z}_{t}\in\mathbb{R}^{n_{s}\times K}$ is the embedding matrix whose $k$th column is $\bm{z}_{t}^{(k)}$, and $\text{vec}(\cdot)$ is the vectorization operator. We refer to the ESN with areal random representation at the input stage as AESN.
3.2 Uncertainty Quantification
The uncertainty quantification for the proposed model follows the ensemble method outlined in McDermott and Wikle, (2017, 2019), incorporating the calibration techniques demonstrated in Daw and Wikle, (2023), Yoo and Wikle, (2023), and Bonas et al., (2024). Specifically, we repeatedly fit the proposed model to the dataset using different sets of randomly sampled weights and potentially varying hyperparameters. For each forecast time $t$ at location $s$, we generate $K$ ensemble forecasts, denoted as $\widehat{Y}^{1}_{t,s},\dots,\widehat{Y}^{K}_{t,s}$. The calibrated prediction interval $(\widehat{l}_{t,s},\widehat{u}_{t,s})$ is then obtained by solving the following optimization problem
$(\widehat{l}_{t,s},\widehat{u}_{t,s})=\operatorname*{argmin}_{(l_{t,s},u_{t,s})}(u_{t,s}-l_{t,s}),$
@@ 3.2 Uncertainty Quantification
where $\alpha$ is the desired significance level. We note that the prediction interval obtained from this approach corresponds to the unimodal highest density region proposed by Hyndman, (1996).
3.3 Hyperparameter selection
Hyperparameter selection plays a critical role in the performance of an ESN, as it significantly impacts the flexibility, stability, and generalization of the model. The most common approaches for hyperparameter selection are grid search (McDermott and Wikle,, 2017, 2019; Lukoševičius,, 2012) and random search (Yoo and Wikle,, 2023; Bergstra and Bengio,, 2012). Another approach involves combining results from ESNs with different hyperparameter sets, which can be viewed as model averaging (Grieshop and Wikle,, 2024). In this work, we opted for the random search method, as it is more efficient in exploring the possible hyperparameter space.
To tune the AESN, we first determine the distribution and its sparsity parameter, following the guidelines of Lukoševičius, (2012) and McDermott and Wikle, (2017). Specifically, we sample $\bm{W}_{\text{res}}$ from a mixture of the Dirac delta function $\delta_{0}$ and a uniform distribution, setting $a_{\text{res}}$ to 0.1. In accordance with Lukoševičius, (2012), we make $\bm{W}_{\text{in}}$ and $\bm{U}^{(k)}$ dense by sampling from a uniform distribution. We then apply random search to tune $a_{u}$, $a_{\text{in}}$, $a_{h}$, $\nu$, $n_{h}$, $k$, and the penalty parameter $\tau$ in ridge regression, by sampling from a log-uniform distribution to find the optimal parameters (Bergstra and Bengio,, 2012).
@@ 4 Tourism Occupancy Analysis
We demonstrate the performance of the proposed method on Eurostat’s tourism occupancy data. Tourism plays a crucial role in shaping regional economies. For instance, infrastructure developed for tourism purposes can significantly boost regional development. Additionally, the creation of numerous job opportunities associated with tourism can enhance the living standards of local residents and contribute to the improvement of local welfare. On the other hand, tourism can pose several challenges to local communities. A large flow of tourists can place considerable pressure on local ecosystems, resulting in significant impact local ecology. Without proper management, increased waste, noise, and transportation emissions can lead to serious pollution issues that threaten local wildlife and contribute to habitat loss. Furthermore, the strain on natural resources and public infrastructure, along with overcrowding, can disrupt residents’ daily lives and place additional stress on local services. These challenges can be mitigated through careful planning and sustainable management practices that leverage tourism occupancy forecasts to minimize negative impacts while maximizing the benefits of tourism (Seghir et al.,, 2015; Brida et al.,, 2016; e Silva et al.,, 2018; Commission and Eurostat,, 2023).
@@ 4.1 Tourism Occupancy Data
The variable used to represent tourism occupancy is the number of nights spent at tourist accommodation establishments, sourced from Eurostat (Eurostat,, 2024; Commission and Eurostat,, 2023). This data is collected through sample surveys or census of various accommodation establishments from business register, and then processed by the competent national statistics authorities (Eurostat,, 2024; Commission and Eurostat,, 2023). In compliance with Commission Delegated Regulation (EU) 2019/1681 of 1 August 2019, these establishments are required to report monthly data on nights spent at tourist accommodations for NUTS  [^5] Nomenclature of Territorial Units for Statistics (NUTS) is the official classification of subregions used by Eurostat. level 2 regions (Commission and Eurostat,, 2023). As defined by the Statistical Classification of Economic Activities in the European Community, these establishments include hotels, holiday and short-stay accommodations, as well as camping grounds, recreational vehicle parks, and trailer parks.
The dataset consists of monthly nights spent at both domestic and foreign tourist accommodations from January 2020 to December 2023, covering 242 NUTS level 2 regions in the EU. We excluded regions with missing data at any time point, resulting in a final dataset of 228 locations with 48 time points each. As shown in Figure 1, due to the geographic nature of EU regions, many isolated areas are present, which can be effectively managed by the self-loops in the areal random representation, as discussed in Section 2.2. For this study, we use the first 36 months as training data and forecast the last 12 months. The night spent are modeled in thousands on a logarithmic scale and transformed back after forecasting.
@@ 4.2 Implementation
For ESN-based models, including standard ESN, ESN with EOF, and AESN, we used a validation set approach to select the optimal hyperparameters. Specifically, we further split the training data into 24 months for training and the remaining 12 months as the validation set. We then performed random searches to identify the best-performing hyperparameters on the validation set, as detailed in Section 3.3. Once the optimal parameters were selected, we refit the ESN based models using the entire 36 months in the training data. For comparison, we used the dlm package in R (Petris,, 2010; Petris et al.,, 2009). The DLM we used in this study included a second-order polynomial component and a seasonal component with a frequency of 12. The parameters in the state and data models were estimated using the maximum likelihood method provided by the dlm package.

### v2

@@ 3.1 Areal Random Representation
In addition to the reservoir, the input signal is another important component of the ESN. It can include covariates that account for confounding effects not captured by the time series alone. When additional covariate variables are unavailable, lagged values are typically used as input signals, as their trajectories describe the temporal dynamics of the series (mcdermott2017ensemble; mcdermott2019deep; yoo2023using). In spatial statistics, spatial input embedding is widely used to capture the spatial structure found in the data. For example, daw2023reds use random Fourier features to approximate the stationary Gaussian process kernel. However, there is no inherent way of incorporating neighborhood structures into the ESN. The common approach for capturing spatial features in ESNs involves modeling the EOFs coefficients asscociated with a spatio-temporal series (mcdermott2017ensemble; mcdermott2019deep; yoo2023using). This approach, however, requires deciding how many EOFs to include and may lead to a potential loss in forecasting performance, as demonstrated in our data analysis.
As mentioned in Section 1, we leverage random representations to capture spatial dynamics at the input stage (bengio2013representation; gallant1987random; sutton2014online). Building on the approach of gallant1987random, we represent spatial features found in areal data through a dense embedding of input signals, using a structure inspired by SGC layers (wu2019simplifying). To motivate our embedding method, we can revisit (3) in the SGC layer. Let $\bm{X}\in\mathbb{R}^{n_{s}\times n_{x}}$ denote the input signals, $\bm{A}\in\mathbb{R}^{n_{s}\times n_{s}}$ the proximity matrix, and $\widetilde{\bm{A}}=\bm{I}+\bm{A}$ the proximity matrix with self-loops. We further let $\widetilde{\bm{D}}$ represent the degree matrix of $\widetilde{\bm{A}}$, and $\bm{S}=\widetilde{\bm{D}}^{-1/2}\widetilde{\bm{A}}\widetilde{\bm{D}}^{-1/2}$ the symmetric normalized adjacency matrix. With some simple linear algebra, (3) can be rewritten as
$\bm{z}^{(k)}(l)=\sum_{l^{\prime}\in N(l)}S(l,l^{\prime})\bm{X}(l)\bm{W}^{(k)},$
where $S(l,l^{\prime})$ denotes the entry of $\bm{S}$ at the $l$-th row and $l^{\prime}$-th column, $\bm{z}^{(k)}(l)$ is the $k$-th embeded feature at location $l$, and $\bm{W}^{(k)}$ is the $k$-th kernel of dimension $n_{x}$. We note that $\bm{W}^{(k)}$ is shared across all locations, which may be less effective at capturing graph features (liu2020comprehensive). To enhance its capacity, we construct the random representation using spatially varying kernels, which we refer to as the areal random representation. The areal random representation is defined as
$\bm{z}^{(k)}(l)$ $=$ $\sum_{l^{\prime}\in\mathcal{N}(l)}S(l,l^{\prime})\bm{X}(l)\bm{U}^{(k)}(l^{\prime}),$ (6)
@@
where $\mathcal{N}(l)$ represents the neighborhood of location $l$ defined by $\widetilde{\bm{A}}$. For $k=1,\dots,K$, $\bm{z}^{(k)}(l)$ is the $k$-th embeded feature at location $l$, and $\bm{U}^{(k)}(l^{\prime})$ are $n_{x}$ randomly sampled weights at location $l^{\prime}$. To further expand the capacity of representation, we generate $K$ different copies of locally averaged random signals for each $\bm{U}^{(k)}$. The resulting embedding matrix $\mathbf{Z}$ is the collection of $\{\mathbf{z}^{(1)},\dots,\mathbf{z}^{(K)}\}$, with dimensions $n_{s}\times K$. Here, we note that the parameter $K$ closely resembles the number of neurons or feature maps in a neural network layer. These weights can also be interpreted as the approximated localized spectral filters with different kernels, enabling the model to capture a rich set of local features. Hence, we embed the input signals at each location from dimension $n_{x}$ to $K$. We also note that these weights are sampled without access to the data, which prevents the embedding layers from learning noise. To control overfitting for the entire AESN model, regularization should be applied to the ridge regression at the output layer. Lastly, we note that (6) can be vectorized as
$\bm{z}^{(k)}=\bm{S}\sum_{j=1}^{n_{x}}(\bm{U}^{(k)}\odot\bm{X})_{j},$ (7)
where $\odot$ is the the Hadamard product. To simplify notation, we use $(\cdots)_{j}$ to denote the $j$-th column of the matrix within the parentheses. In (7), we first compute the column sums of $(\bm{U}^{(k)}\odot\bm{X})$, then apply a matrix product with $\bm{S}$ to filter out relevant neighbors.
To integrate these embedded signals into the ESN, we apply the areal random representation to the input signals $\widetilde{\bm{x}}_{t}$ at each time step, resulting in the embedded signals $\bm{z}_{t}^{(k)}$, as shown in (11). The vectorized signals $\widetilde{\bm{z}}_{t}$ are then used as the input to the ESN, as in (10). This implies that when a signal at any location is activated in the reservoir stage, all signals in its neighborhood are also activated. AESN can be defined as,
@@
$\bm{h}_{t}$ $=(1-\alpha)\bm{h}_{t-1}+\alpha\widetilde{\bm{h}}_{t}$ (8)
$\widetilde{\bm{h}}_{t}$ $=g_{h}(\frac{\nu}{|\lambda_{w}|}\bm{W}_{res}\bm{h}_{t-1}+\bm{W}_{in}\widetilde{\bm{z}}_{t})$ (9)
@@
$\bm{z}_{t}^{(k)}$ $=\bm{S}\sum_{j=1}^{n_{x}}(\bm{U}^{(k)}\odot\widetilde{\bm{x}}_{t})_{j}$ (11)
$\bm{U}^{(k)}$ $\sim\text{Unif}(-a_{u},a_{u}),$ (12)
where $\bm{Z}_{t}\in\mathbb{R}^{n_{s}\times K}$ is the embedding matrix whose $k$th column is $\bm{z}_{t}^{(k)}$, and $\text{vec}(\cdot)$ is the vectorization operator. Here, it is worth emphasizing that capturing spatial dynamics at the input stage offers two key advantages. First, manipulating the input stage preserves the all desirable properties of the ESN, such as the ESP. Second, handling spatial dynamics at the input stage introduces minimal computational overhead, as it avoids being recursively computed at every time step.
3.2 Hyperparameter selection
Hyperparameter selection plays a critical role in the performance of an ESN, as it significantly impacts the flexibility, stability, and generalization of the model. The most common approaches for hyperparameter selection are grid search (mcdermott2017ensemble; mcdermott2019deep; lukovsevivcius2012practical) and random search (yoo2023using; bergstra2012random). Another approach involves combining results from ESNs with different hyperparameter sets, which can be viewed as model averaging (grieshop2024echo). In this work, we opted for the random search method, as it is more efficient in exploring the possible hyperparameter space.
To tune the AESN, we first determine the distribution and its sparsity parameter, following the guidelines of lukovsevivcius2012practical and mcdermott2017ensemble. Specifically, we sample $\bm{W}_{\text{res}}$ from a mixture of the Dirac delta function $\delta_{0}$ and a uniform distribution, setting $a_{\text{in}}$$a_{\text{res}}$} to 0.1. In accordance with lukovsevivcius2012practical, we make $\bm{W}_{\text{in}}$ and $\bm{U}^{(k)}$ dense by sampling from a uniform distribution. We then apply random search to tune $a_{u}$, $\nu$, $n_{h}$, $k$, and the penalty parameter $\tau$ in ridge regression. Specifically, we split the validation data from the training data. In each iteration of the random search, we sample the hyperparameters from log-uniform distributions. The search ranges are $(15,300)$ for $K$ and $n_{h}$, $(0.1,1)$ for $a_{u}$, $(0.05,1)$ for $\nu$, and $(0.001,0.1)$ for $\tau$. Here, the log-uniform distribution is used for its efficiency in exploring values that vary across different orders of magnitude (bengio2013representation). For integer parameters, we round the sampled values to the nearest integer. The optimal set of parameters is selected based on the best forecast performance on the validation set.
3.3 Ensemble AESN
The uncertainty quantification for the proposed model follows the ensemble method outlined in mcdermott2017ensemble; mcdermott2019deep, incorporating the calibration techniques demonstrated in daw2023reds, yoo2023using, and bonas2024calibrated. Specifically, we repeatedly fit the proposed model to the dataset using different sets of randomly sampled weights and potentially varying hyperparameters. For each forecast time $t$ at location $s$, we generate $K$ ensemble forecasts, denoted as $\widehat{Y}^{1}_{t,s},\dots,\widehat{Y}^{K}_{t,s}$. The calibrated prediction interval $(\widehat{l}_{t,s},\widehat{u}_{t,s})$ is then obtained by solving the following optimization problem
$(\widehat{l}_{t,s},\widehat{u}_{t,s})=\operatorname*{argmin}_{(l_{t,s},u_{t,s})}(u_{t,s}-l_{t,s}),$ (13)
@@ 3.3 Ensemble AESN
where $\alpha$ is the desired significance level. We note that the prediction interval obtained from this approach corresponds to the unimodal highest density region proposed by hyndman1996computing. While we acknowledge that the behavior of ESN forecasts can be quite complex, we choose to use the unimodal highest density region for computational efficiency, as commonly demonstrated in the ESN literature (yoo2023using; daw2023reds; bonas2024calibrated). Although this unimodal interval is preferred for practical reasons, users may still apply any calibration method after generating the ensemble of forecasts.
Finally, we include the full algorithm for ensemble AESN Algorithm 1.
Algorithm 1 Ensemble Forecasting with AESN
1:Input: $\{\bm{Y}_{t}\}$ and $\{\tilde{\bm{x}}_{t}\}$ for $t=1,\dots,T$
2:Tuning: Select optimal parameters $\{k,a_{u},a_{in},a_{res},\gamma^{in},\gamma^{res},\alpha,\nu\}$ using random search and empirical knowledge.
3:for $l=1,\dots,L$ do
4:  Simulate $U$ using (12)
5:  Calculate $\{z_{t}:t=1:T\}$ using (11)
6:  Simulate $W_{res}$ and $W_{in}$ using (1) and (2)
7:  Calculate $\{h_{t}:t=1:T\}$ using (8) and (9)
8:  Estimate $W_{out}$ using ridge regression
9:  Obtain forecasts $\widehat{\bm{Y}}_{T+1}^{l}$
10:end for
11:Denote the ensemble of forecasts $\widehat{\bm{Y}}_{T+1}^{ensemble}=\{\widehat{\bm{Y}}_{T+1}^{l}:l=1,\dots L\}$
12:Output ensemble forecast mean $\overline{\bm{Y}}_{T+1}=\text{mean}(\widehat{\bm{Y}}_{T+1}^{ensemble})$, ensemble forecast interval $(\widehat{l}_{t,s},\widehat{u}_{t,s})$ according to (13)
@@ 4 Tourism Occupancy Analysis
We demonstrate the performance of the proposed method on Eurostat’s tourism occupancy data. Tourism plays a crucial role in shaping regional economies. For instance, infrastructure developed for tourism purposes can significantly boost regional development. Additionally, the creation of numerous job opportunities associated with tourism can enhance the living standards of local residents and contribute to the improvement of local welfare. On the other hand, tourism can pose several challenges to local communities. A large flow of tourists can place considerable pressure on local ecosystems, resulting in significant impacts on local ecology. Without proper management, increased waste, noise, and transportation emissions can lead to serious pollution issues that threaten local wildlife and contribute to habitat loss. Furthermore, the strain on natural resources and public infrastructure, along with overcrowding, can disrupt residents’ daily lives and place additional stress on local services. These challenges can be mitigated through careful planning and sustainable management practices that leverage tourism occupancy forecasts to minimize negative impacts while maximizing the benefits of tourism (seghir2015tourism; brida2016has; e2018analysing; doi/10.2785/606702).
@@ 4.1 Tourism Occupancy Data
The variable used to represent tourism occupancy is the number of nights spent at tourist accommodation establishments, sourced from Eurostat (tour-occ-nin2m-2024-9-12; doi/10.2785/606702). This data is collected through sample surveys or census of various accommodation establishments from business register, and then processed by the competent national statistics authorities (tour-occ-nin2m-2024-9-12; doi/10.2785/606702). In compliance with Commission Delegated Regulation (EU) 2019/1681 of 1 August 2019, these establishments are required to report monthly data on nights spent at tourist accommodations for NUTS level 2 regions (doi/10.2785/606702). As defined by the Statistical Classification of Economic Activities in the European Community, these establishments include hotels, holiday and short-stay accommodations, as well as camping grounds, recreational vehicle parks, and trailer parks.
The dataset consists of monthly nights spent at both domestic and foreign tourist accommodations from January 2020 to December 2023, covering 242 NUTS level 2 regions in the EU. We excluded regions with missing data at any time point, resulting in a final dataset of 228 locations with 48 time points each. As shown in Figure 1, due to the geographic nature of EU regions, many isolated areas are present, which can be effectively managed by the self-loops in the areal random representation, as discussed in Section 2.2. For this study, we use the first 36 months as training data and forecast the last 12 months. In addition, we use the lagged values of the inputs, with the optimal number of lagged values determined together with the hyperparameters. In this analysis, the optimal number of lagged values is 5. The night spent are modeled in thousands on a logarithmic scale and transformed back after forecasting.
@@ 4.2 Implementation
For ESN-based models, including standard ESN, ESN with EOF, and AESN, we used a validation set approach to select the optimal hyperparameters. Specifically, we further split the training data into 24 months for training and the remaining 12 months as the validation set. We then performed random searches to identify the best-performing hyperparameters on the validation set, as detailed in Section 3.2. Once the optimal parameters were selected, we refit the ESN based models using the entire 36 months in the training data. We used 500 ensembles for all ESN-based models, as this ensures consistent results for our data analysis. We present the relevant optimal hyperparameters for AESN in Table 1. Sensitivity analysis shows that AESN tends to favor smaller values of $K$, $n_{h}$, and $a_{u}$ for simpler tasks such as the 3-step forecast, while it prefers larger values for more challenging tasks like the 12-step forecast. We also note that due to its randomness, the random search may not always find the most parsimonious configurations. For the graph construction, we use unweighted graph with the Queen adjacency matrix, where diagonal entries are set to zero and non-zero off-diagonal entries indicate neighboring relationships. In practice, users may also consider alternative weighted graphs, such as those based on economic distance. The selection of the number of EOFs follows a randomization approach. Specifically, we compare the singular values of the training data with those of an equivalent uncorrelated dataset, created by randomly rearranging each column of the training data. The optimal number of EOFs is chosen as the point where the traces of the singular values intersect (hastie2009elements; wikle2019spatio). Using this method, the first 8 EOFs are selected. For comparison, we used the dlm package (dlm_r; dlm_book). The DLM we used in this study included a second-order polynomial component and a seasonal component with a frequency of 12. The parameters in the state and data models were estimated using the maximum likelihood method provided by the dlm package. We implemented all four models on an AMD Ryzen 5 7640U (12) @ 4.97 GHz CPU with 32 GB of RAM. The DLM model was implemented using the dlm package in R, and all ESN-based models were implemented using NumPy in Python (harris2020array).
Table 1: Optimal Parameters for AESN
Forecast $K$ $a_{u}$ $n_{h}$ $\nu$
3-step 16 0.094 168 0.414
12-step 122 0.958 253 0.072